Find an equation of the line containing points(7,9) and parallel to the line x+9y=2..
Answer by ewatrrr(24785)


Hi,
using the general slope-intercept form for an equation of a line
y= mx + b where m is the slope and b the y intercept
line parallel to the line x+9y= 2
x+9y = 2
y = -x/9 + 2/9 New line will have slope (-1/9)
y = (-1/9)x + b Using Pt(7,9) to solve for b
9 = (-1/9)*7 + b
81/9 + 7/9 = b
88/9 = b
y = (-1/9)*7 + 88/9